OpenStudy (anonymous):

A trapezoid has an area of 22 square inches. The longer base measures 6 inches and the height is 4 inches. use the formula A=1/2h(b1+b2) to find the length of the shorter base. a. b = 10 b. b = 5 c. b = 7.5 d. b = 12.5

OpenStudy (ltrout):

The easiest way to do this would be to fill in the equation with each answer choice. \[1) A = 1/2(4)(6 + 10)\]\[2) A = 1/2(4)(6+5)\]\[3) A = 1/2(4)(6+7.5)\]\[4) A = 1/2(4)(6+12.5)\] And solve each of those until you get the one that makes a = 22 (since we already know what the area is.

OpenStudy (tkhunny):

A=1/2h(b1+b2) Or, of course, one could use one's best algebra skills and solve directly. 22=(1/2)4(6+b2) Multiply by 2 Divide by 4 Etc. However, we are told that 6 is the longer base. How many of those answer choices are shorter than 6"? It's an eyeball problem. Pick Short Base = 5" and be done!

OpenStudy (anonymous):

so number 1 then right????

OpenStudy (ltrout):

No. Fill in that equation I gave you at the top and see if number one equals out to 22.

OpenStudy (tkhunny):

You're not paying attention. You didn't read what I wrote. "longer base measures 6 inches " Is 10" shorter than 6"?

OpenStudy (anonymous):

no its not

OpenStudy (tkhunny):

Is 5" shorter than 6"

OpenStudy (anonymous):

yes

OpenStudy (tkhunny):

Is there any choice besides 5" that is less than 6"?

OpenStudy (anonymous):

no